OpenStudy (anonymous):

Find the first integer \(n\ge 1\) such that the average of \(1^2;2^2;3^2...n^2\) is itself a perfect square.

OpenStudy (mathmale):

Keep in mind that the "average" of a set of numbers is \[average=\frac{ \sum~of~the~numbers }{ count~of~the~numbers }\]

OpenStudy (mathmale):

for example, the average of 1^2, 2^2 and 3^2 is\[average = \frac{ 1^2+2^2+3^2 }{ 3 }=\frac{ 1+4+9 }{ 3 }=\frac{ 14 }{ 3 }\] Is (14/3) a perfect square? No? Then you'll have to keep going; next would be the average of the squares of 1, 2, 3 and 4. I don't know of a more sophisticated approach; perhaps someone else does. Nevertheless, by continuing to find the averages as explained here, you'll likely eventually find the desired integer n.

ganeshie8 (ganeshie8):

the trivial answer is n = 1 for other solutions, we may try to mess wid : \(\large \frac{(n+1)(2n+1)}{6} = k^2\) \(\large (n+1)(2n+1) = 6k^2\)

ganeshie8 (ganeshie8):

i feel stuck... @KingGeorge

OpenStudy (mathmale):

Hey, ganeshie8, that's a great approach!! I'd suggest you look up "summation formula n^2" or something equivalent. There is definitely a formula for the sum of the terms k^2, but I've forgotten it. Once we have that formula, it'd be an easy matter to let n = 2, 3, 4, ..., until we find a sum that is itself a perfect square. More power to you!

OpenStudy (mathmale):

See http://polysum.tripod.com/ You'll find the necessary summation formula for k^2 there.

ganeshie8 (ganeshie8):

yeah looks we're not going to get lucky wid small numbers the first solution after n = 1 is n= 337 which is a very messy prime lol

OpenStudy (mathmale):

How'd you arrive at n=337?

OpenStudy (mathmale):

For the lack of a more sophisticated approach, I'd set up an Excel spreadsheet showing the sums of k^2 for n=1, 2, 3, etc. Grunt work. But it'd surely show us quickly for which n the sum of k^2 would itself be a perfect square.

OpenStudy (anonymous):

can we use some elementary number theory to find the roots,like this \(6|(n+1)(2n+1)\)

OpenStudy (anonymous):

trial and error

ganeshie8 (ganeshie8):

lets see... 2n + 1 is odd so 2 | (n+1)

OpenStudy (anonymous):

which implies that n is itself odd,otherwise that will not hold

ganeshie8 (ganeshie8):

yahhh thats pretty much it, we cannot figure out anything further from that i think...

OpenStudy (anonymous):

hmmm.....ok ,as alternative,is it not enough for us to seek to write \(\dfrac{(n+1)(2n+1)}{6}\) in the form \((an+b)^2\) for \(a,b\) real numbers

OpenStudy (anonymous):

this doesnt seem to be possible :/

OpenStudy (anonymous):

another try at this....it seems like if we try to solve this using quadratic equation we have \[n=\dfrac{-3+\sqrt{1+48k^2}}{2(2)}\implies \] \[1+48k^2\] hould also be a square

ganeshie8 (ganeshie8):

(4n+3)^2 = 48k^2 + 1

OpenStudy (anonymous):

n=1,k=1 so we also remember that n is odd so \[4n+3=4(2j-1)+3=(8j-1)\] \[(8j-1)^2=48k^2+1\implies 8j(8j-2)=48k^2\] \[j(4j-1)=3k^2\]

OpenStudy (anonymous):

is this taking us to square zero,i mean back to our original problem? or leading somewher

OpenStudy (anonymous):

i'm taking it to SE,i will update it as it gets solutions there :)

ganeshie8 (ganeshie8):

Ohhk... good luck :) plz provide the SE link also.....

OpenStudy (anonymous):

thanks a lot ,its nice brainstorming with you,u have great approach to problem solving Sir

ganeshie8 (ganeshie8):

lol dont make me Sir xo Jonask ever tried quadratic diaphantine equations ?

OpenStudy (anonymous):

yes i have tried them hmmmm ok now i have a new approach all together,but will need some help

OpenStudy (anonymous):

i only studied diophantine of the form \[ax^2+bxy+cy^2=m\]

ganeshie8 (ganeshie8):

I have been lazy to complete the last 4 chapters in NT.. which cover all these cool stuff :/ looks the equation u got wid quadratic formula can be thought of as a quadratic : \((8j-1)^2=48k^2+1\) \((8j-1)^2-48k^2+1\)

ganeshie8 (ganeshie8):

*\((8j-1)^2-48k^2 = 1 \)

ganeshie8 (ganeshie8):

i think it should be easy for u to solve the quadratic diaphantine.... for me it wud take some time to udnerstand the mechanics of quadratic equations....

OpenStudy (anonymous):

this is perfect,now we have found an organized solution,actually after u sujjested this,i think i dont need SE

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/DiophantineEquation2ndPowers.html x^2 - Dy^2 = 1

OpenStudy (anonymous):

i will take some time as well,i'm sure wolfram can do a better job

ganeshie8 (ganeshie8):

wow ! plz show the solution once u have it ! would love to see how to get n=337 analytically xD